OpenStudy (anonymous):

Write the equation of the like passing through the points (5,-7) and (-4,2).

OpenStudy (anonymous):

(y0-y1)=m(x0-x1)

OpenStudy (anonymous):

@hannahellens221 Find slope. Then use the Point-Slope formula like @liliy said. (: Heres an example, if it helps. http://openstudy.com/study#/updates/50b2d27ce4b0e906b4a65d89

OpenStudy (anonymous):

change in y over change of x gives you the slope: y+y /x+X: -5/1..m=-5

OpenStudy (anonymous):

thank you so much!!!!